OpenStudy (anonymous):

summation from n=2 to infinity (n^.5 +4)/n^2. Prove it's convergent by comparison.

OpenStudy (amistre64):

for large values of n, this is pretty much sqrt(n)/n^2, or 1/n^(3/2)

OpenStudy (amistre64):

\[ \frac{n^.5 +4}{n^2}<\frac{1}{n}^{(b)}\] \[ n^b<\frac{n^2}{n^.5 +4}\] \[ ln(n^b)<ln(\frac{n^2}{n^.5 +4})\] \[ b~ln(n)<ln(n^2)-ln({n^.5 +4})\] \[ b<\frac{ln(n^2)}{ln(n)}-\frac{ln({n^.5 +4})}{ln(n)}\] \[ b<\lim_{n\to inf}\frac{ln(n^2)}{ln(n)}-\frac{ln({n^.5 +4})}{ln(n)}\]

OpenStudy (amistre64):

\[\lim_{n\to inf}\frac{ln(n^2)}{ln(n)}-\frac{ln({n^.5 +4})}{ln(n)}\] \[\lim_{n\to inf}1+ln(n)-\frac{ln({n^.5 +4})}{ln(n)}\] after a long and arduous process, b <= 3/2 should suffice